OpenStudy (anonymous):

Which of the following is a solution of x2 + 6x = −18? x = 3 − 3i x = −3 + 3i x = −6 + 3i x = 6 − 3i

OpenStudy (anonymous):

it will bbecome x^2+6x+18 =>now use quadratic formua .. x=-6+3i

OpenStudy (anonymous):

now for this were completing the square..

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

yuppp...complete it like x2+2(x)(3)+(3)^2-(3)^2+18 =>{(x+3)^2}-9+18 =?(x+3)^2+9

OpenStudy (anonymous):

I did it like this x^2+6x=-18 6/2=3^2=9 x^2+6x+9=-18+9 x^2+6x+9=-9

OpenStudy (anonymous):

@zaibali.qasmi

OpenStudy (anonymous):

Use this formula and you will get it http://www.mathsisfun.com/algebra/quadratic-equation-derivation.html

OpenStudy (anonymous):

And you know that \[\sqrt{-1}= i\]

OpenStudy (anonymous):

yuppp..its too correct..

OpenStudy (anonymous):

@brijackson6